# Iterative assembly of ${}^{171}$b atom arrays in cavity-enhanced optical lattices

### v2

@@ VI Appendix B: Sources of Per-Cycle Loss
Figure 5: Characterization of per-cycle loss mechanisms. (a), (b). Loss per pair of handoffs between the 459 nm (a) and 423 nm (b) tweezer arrays and lattice, measured by performing 25 handoffs with cooling in between. The black line represents the estimated contribution from vacuum loss, with the red band representing day-to-day drifts in this value. Tweezer depths are normalized to their default values, which correspond to roughly 50 $\mu$K. (c) Final vacancy fraction versus rearrangment tweezer power, normalized to the default tweezer depth of approximately 150 $\mu$K. For low powers, atoms are not efficiently transferred from the reservoir to target array. For too high of powers, we observe increased loss of loaded atoms due to the rearrangement tweezer passing nearby.
@@
Vacuum loss due to the finite duration of loading cycles is the single largest source of loss in our system, contributing percent-level per-cycle loss. Because our loading cycle is of variable length, and because the vacuum level in our system can fluctuate from day to day, it is difficult to estimate its exact contribution. However, typical values for the loading cycle duration and vacuum lifetime are 300 ms and 30 s respectively, so 1% represents a typical value for this loss. In this section, we describe other loss sources that contribute to a lesser degree, especially if parameters are not carefully optimized. In general, the loss mechanisms described below both limit the size of array we can ultimately load (as the reservoir must be large enough to replace lost atoms), and limit the final fill fraction of the array.
@@
As described in the main text, typical imaging loss is $2(1)\times 10^{-3}$ where roughly 30 % is the vacuum loss and the rest is the Raman scattering out of ${}^{3}$${}_{1}$ due to the trap light , and the discrimination infidelity is typically at or below the $10^{-3}$ level. If an atom within the target array is lost during the image after scattering enough photons to be identified as present, the defect will not be filled during the subsequent rearrangement step, leading to a defect in the array. If a site within the array is incorrectly identified as empty, an atom will likely be added to that site from the reservoir, and subsequently undergo lossy collisions with the original occupant. Empty sites within the target or reservoir arrays that are mistakenly identified as full will lead to a previous defect not being repaired. However, because defects are rare, this mechanism is less problematic. We currently operate with a discrimination threshold that balances the correct identification of empty and full sites. In a system where imaging losses and infidelity become dominant loss sources (a system with better vacuum), it may be advantageous to bias the threshold to minimize the more problematic forms of imaging error.
@@
Each image requires handing off atoms from the tweezer arrays into the lattice and back. We find that this process is sensitive to the alignment of the two arrays on the scale of a single lattice site, to the depth of the tweezer arrays, and to the temperature of the atoms. Atom temperature and alignment are described in Appendix A and C, respectively. We isolate the effect of trap depth on handoff for cold atoms and well-aligned arrays by performing 25 subsequent handoff pairs between tweezers and lattice in figure 5a, b. We perform our usual Raman sideband cooling (RSC) each time the atoms are in the lattice between handoffs. For 459 nm tweezer depths below 75% of our typical operating conditions, we observe significant loss. Above this power level, we observe a constant loss rate of 0.0006(1), which is consistent with data taken with the handoff omitted, and with our typical vacuum losses. The 423 nm tweezers show a similar behavior, though with a slightly higher loss rate of 0.0015(2) at high powers, likely due to worse alignment or optical aberrations. If the tweezers are aligned to the trough of the XY cavity intensity, the loss can be as high as 0.02.
@@
In principle, moving new atoms in the target array may cause loss of existing atoms in the array, as the rearrangement tweezer moves near the occupied target sites. We assess this possibility in figure 5c by measuring the final fill fraction versus the depth of the rearrangement tweezer. For a rearrangement tweezer that is too shallow, we observe a lower loading rate and so a lower final fill fraction. For too deep a tweezer, we also see a reduction in the final fill fraction, as the rearrangement leads to loss of existing atoms. Near our typical operating conditions, we observe a region with weak dependency of the fill fraction on the tweezer power.
VII Appendix C: Alignment of arrays
@@
We then optimize the translation alignment by performing repeated handoffs between the tweezers and the lattice and scanning the position offset of the tweezer arrays. We observe a periodic modulation in the atomic survival (fig. 6), which we fit to determine the optimal alignment. This process is repeated as necessary to ensure alignment of the arrays.

### v3

@@ Appendix F Sources of Per-Cycle Loss
Figure 6: Characterization of per-cycle loss mechanisms. (a), (b). Loss per pair of handoffs between the 459 nm (a) and 423 nm (b) tweezer arrays and lattice, measured by performing 25 handoffs with cooling in between. The black line represents the estimated contribution from vacuum loss, with the red band representing day-to-day drifts in this value. Tweezer depths are normalized to their default values, which correspond to roughly 50 $\mu$K. The higher loss present in the 423 nm tweezer array is likely due to higher levels of optical aberrations compared to the 459 nm traps. (c) Final vacancy fraction versus rearrangement tweezer power, normalized to the default tweezer depth of approximately 150 $\mu$K. For low powers, atoms are not efficiently transferred from the reservoir to target array. For too high of powers, we observe increased loss of loaded atoms due to the rearrangement tweezer passing nearby.
@@
Background gas collisions during the finite duration of loading cycles is the single largest source of loss in our system, contributing percent-level per-cycle loss. Because our loading cycle is of variable length, and because the vacuum level in our system can fluctuate from day to day, it is difficult to estimate its exact contribution. However, typical values for the loading cycle duration and vacuum lifetime are 300 ms and 30 s respectively, so 1% represents a typical value for this loss. In this section, we describe other loss sources that contribute to a lesser degree, especially if parameters are not carefully optimized. In general, the loss mechanisms described below both limit the size of array we can ultimately load (as the reservoir must be large enough to replace lost atoms), and limit the final fill fraction of the array.
@@
As described in the main text, typical imaging loss is $2(1)\times 10^{-3}$ where roughly 30 % is the vacuum loss and the rest is the Raman scattering out of 3P1 due to the trap light, and the discrimination infidelity is typically at or below the $10^{-3}$ level (see appendix G). The contribution from Raman scattering could be mitigated by applying repump lasers to the 3P0 and 3P2 states, though spin selectivity would be compromised in applications where that is required. If an atom within the target array is lost during the image after scattering enough photons to be identified as present, the defect will not be filled during the subsequent rearrangement step, leading to a defect in the array. If a site within the array is incorrectly identified as empty, an atom will likely be added to that site from the reservoir, and subsequently undergo lossy collisions with the original occupant. Empty sites within the target or reservoir arrays that are mistakenly identified as full will lead to a previous defect not being repaired. However, because defects are rare, this mechanism is less problematic. We currently operate with a discrimination threshold that balances the correct identification of empty and full sites. In a system where imaging losses and infidelity become dominant loss sources (a system with better vacuum), it may be advantageous to bias the threshold to minimize the more problematic forms of imaging error.
@@
Each image requires handing off atoms from the tweezer arrays into the lattice and back. We find that this process is sensitive to the alignment of the two arrays on the scale of a single lattice site, to the depth of the tweezer arrays, and to the temperature of the atoms. Atom temperature and alignment are described in appendix E and H, respectively. We isolate the effect of trap depth on handoff for cold atoms and well-aligned arrays by performing 25 subsequent handoff pairs between tweezers and lattice in figure 6a, b. We perform our usual Raman sideband cooling (RSC) each time the atoms are in the lattice between handoffs. For 459 nm tweezer depths below 75% of our typical operating conditions, we observe significant loss. Above this power level, we observe a constant loss per handoff pair of 0.0006(1), which is consistent with data taken with the handoff omitted, and with our typical vacuum losses. The 423 nm tweezers show a similar behavior, though with a slightly higher loss per handoff pair of 0.0015(2) at high powers, likely due to worse alignment or optical aberrations. If the tweezers are aligned to the trough of the XY cavity intensity, the loss can be as high as 0.02.
@@
In principle, moving new atoms in the target array may cause loss of existing atoms in the array, as the rearrangement tweezer moves near the occupied target sites. In our rearrangement geometry, the moving tweezer passes 1.6 $\mu$m – half the array spacing – from the target sites. We assess this possibility in figure 6c by measuring the final fill fraction versus the depth of the rearrangement tweezer. For a rearrangement tweezer that is too shallow, we observe a lower loading rate and so a lower final fill fraction. For too deep a tweezer, we also see a reduction in the final fill fraction, as the rearrangement leads to loss of existing atoms. Near our typical operating conditions, we observe a region with weak dependency of the fill fraction on the tweezer power.
Appendix G Determining imaging fidelity
We use several methods to estimate our imaging fidelity. The simplest (suitable for estimating the fidelity from limited data) is to generate a histogram of the number of collected photons (as shown in figure 3b) and fit a sum of two Gaussians to the peaks corresponding to background counts (indicating no atom in the imaged state) and signal counts (indicating an atom in the imaged state). A threshold can then be defined to distinguish the presence or absence of an atom in the imaged state, and the fidelity estimated by the normalized area of the two Gaussians on the incorrect side of the threshold. From this method, we estimate a typical imaging fidelity of 99.95%. While this method is convenient as it allows the fidelity to be estimated from datasets used for other purposes, it is not rigorous because the signal and background peaks may not be well-modeled by Gaussians.
To provide a more rigorous calibration of the imaging fidelity, we perform many images in a row. We identify misidentification events as instances when the qubit state identified in an image differs from the state in both the images that precede and follow it. We perform this characterization in a stochastically loaded array (50% filling), which evenly samples the possibility of false positive and negatives. We divide the number of misidentification events by the total number of atom images (6570, excluding first and last images, as they cannot meet the criteria for misidentification) to extract a fidelity of $99.92^{+0.03}_{-0.04}$%. For our chosen threshold, the infidelity is evenly balanced between false positives and negatives, to within our statistical uncertainty.
Appendix H Alignment of arrays
@@
We then optimize the translation alignment by performing repeated handoffs between the tweezers and the lattice and scanning the position offset of the tweezer arrays. We observe a periodic modulation in the atomic survival (fig. 7), which we fit to determine the optimal alignment. This process is repeated as necessary to ensure alignment of the arrays (typically once per hour).
Appendix I Comparison of cavity-enhanced lattice to tweezers
Cavity-enhanced optical lattices and optical tweezers represent two alternative methods to generating large number of deep optical traps. Here, we compare the efficiency of the two approaches in the context of generating large arrays of traps for atoms separated by distances appropriate for individual addressing and Rydberg interactions. A third method – using high-power lasers to generate deep lattices without the use of cavity enhancement – can be inferred from the cavity analysis with a build-up factor of 1.
For concreteness, we consider a square two-dimensional grid of $N$ lattice sites, separated by a distance $a$, trapped in either optical lattices or tweezers. In the case of lattices, the grid is located at the intersection of several lattice beams, which can either be configured to interfere with each-other or not. We define the depth $U$ as the amount of energy a classical atom would take to escape its local confinement – the barrier between adjacent sites for the lattice, or the total depth for the tweezers – and neglect gravity because its effect is negligible for traps deep enough to be suitable for rapid imaging. For a cavity-enhanced lattice, the depth of the deepest traps is given by:
$U_{l}=2\alpha P\Lambda I/\pi w^{2}$
Here, $\alpha$ is the atomic polarizability. $P$ is the laser power delivered to the cavity. $\Lambda=2F\epsilon/\pi$ is the power buildup factor. $F$ is the cavity finesse and $\epsilon=T_{in}/(L_{tot}+T_{tot})$ quantifies the impedance matching to the cavity. $T_{in}$ is the transmission of the input mirror, $T_{tot}$ is the total round-trip transmission of all mirrors in the cavity, and $L_{tot}$ is the total additional loss. $w$ is the mode waist of the cavity at the location of the atoms. $I$ quantifies the degree of interference at the location of the atoms. For a simple standing-wave cavity, $I=4$. For a self-intersecting standing-wave cavity like our XY cavity, $I$ = 16. Note that the self-intersecting cavity is a factor of eight more efficient because power must be split between two simple cavities to provide two directions of confinement. A self-intersecting travelling-wave cavity like our Z cavity has $I=4$. The scaling of required power with atom number $N$ comes from the required increase of $w$ with larger atom number.
For tweezers, the trap depth can be approximated as:
$U_{t}=2\alpha P/\pi w_{t}^{2}N$
Here, $P$ is the total power used to form the tweezers, $w_{t}$ is the mode waist of the tweezers (approximated as Gaussian beams).
For our cavity parameters and the atomic polarizability for 783.8 nm ($\alpha=0.79$ MHz/mW/$\mu\rm{m}^{2}$), $U_{l}/P=0.20$ MHz/mW for the XY cavity, and $U_{l}/P=0.10$ MHz/mW for the Z cavity. We have set $\epsilon=0.96$ here to represent the fact that the loss in both cavities is dominated by equal transmission through an input and output mirror (the other two mirrors in each cavity are coated for much higher reflectivity) and the measured bound on per-mirror loss. For comparison, we may consider a tweezer array of 1225 traps with a waist $w_{t}=700~{}nm$, which is typical for tweezer systems operating with wavelengths near 784 nm. In this case, $U_{t}/P=0.00083$ MHz/mW. For this example, our XY cavity requires 255 times less power than the tweezer array to achieve the same depth, while the Z cavity requires 132 times less power. Note that to provide three-dimensional confinement between adjacent sites in the lattice with a given depth, we must provide that depth of trapping with both the XY and Z cavities. We have also neglected the efficiency of delivering light to the cavities relative to tweezers, but in practice the efficiency can be similar for the cavities and tweezers. This dramatic improvement in power efficiency is especially important for architectures that require or benefit from the use of magic-wavelength traps, as this can constrain the use of high-power lasers.